# GRAph-Structured Pyramidal Whole Slide Image Representation

### v1

@@ Abstract
Cancer subtyping is one of the most challenging tasks in digital pathology, where Multiple Instance Learning (MIL) by processing gigapixel whole slide images (WSIs) has been in the spotlight of recent research. However, MIL approaches do not take advantage of inter- and intra-magnification information contained in WSIs. In this work, we present GRASP, a novel graph-structured multi-magnification framework for processing WSIs in digital pathology. Our approach is designed to dynamically emulate the pathologist’s behavior in handling WSIs and benefits from the hierarchical structure of WSIs. GRASP, which introduces a convergence-based node aggregation instead of traditional pooling mechanisms, outperforms state-of-the-art methods over two distinct cancer datasets by a margin of up to 10% balanced accuracy, while being 7 times smaller than the closest-performing state-of-the-art model in terms of the number of parameters. Our results show that GRASP is dynamic in finding and consulting with different magnifications for subtyping cancers and is reliable and stable across different hyperparameters. The model’s behavior has been evaluated by two expert pathologists confirming the interpretability of the model’s dynamic. We also provide a theoretical foundation, along with empirical evidence, for our work, explaining how GRASP interacts with different magnifications and nodes in the graph to make predictions. We believe that the strong characteristics yet simple structure of GRASP will encourage the development of interpretable, structure-based designs for WSI representation in digital pathology. Furthermore, we publish two large graph datasets of rare Ovarian and Bladder cancers to contribute to the field.
@@ 1 Introduction
Figure 1: Representing a multi-magnification WSI via a hierarchical graph can capture inter- and intra-magnification information
@@
Though deep learning has revolutionized computer vision in many fields, digital pathology tasks such as cancer classification remain a complex problem in the domain. For natural images, the task usually relates to assigning a label to an image with an approximate size of 256 × 256 pixels, with the label being clearly visible and well-represented in the image. Gigapixel tissue whole-slide images (WSIs) break this assumption in digital pathology as images exhibit enormous heterogeneity and can be as large as 150,000 × 150,000 pixels. Further, labels are provided at the slide level and may be descriptive of a small region of pixels occupying a minuscule portion of the total image, or they may be descriptive of complex interactions between the substructures within the entire composition of the WSI [9, 28, 20].
@@
Multiple Instance Learning (MIL) has become the prominent approach to address the computational complexity of WSI; however, the majority of methods in the literature focus only on a single level of magnification, usually $20x$ [18, 15, 23, 29, 5, 25, 30, 10]. Using this magnification, a set of patches from each WSI are extracted and used as an instance-level representation. This neither captures the biological structure of the data nor does it follow the diagnostic protocols of pathologists. That is to say, WSIs at higher magnifications contain additional details about the cell nucleus and intra- and extra-cellular matrix, while at lower magnifications one might discern the presence of blood vessels, connective tissue, or muscle fibers. Further, these structures are inconsistent from patient to patient, slide to slide, and subtype to subtype [19]. To capture this variability, pathologists generally use a variety of lenses in their inspection of a tissue sample under the microscope, switching between different magnifications as needed. They generally begin with low magnifications to identify regions of interest for making preliminary decisions before increasing magnifications to confirm or rule out diagnoses [21].
To address this challenge, several multi-magnification approaches have recently been introduced for different tasks of cancer subtyping, survival analysis, and image retrieval, though they are hindered as they often possess millions of parameters [26, 17, 22, 4, 8, 11].
In this research, we aim to further the progress of deep learning in this context by leveraging graph convolutional neural networks (GCNs) [16] with a multi-magnification component as they have been shown to capture more comprehensive and interpretable histopathologic information relating to the underlying mechanisms of a disease [1]. To this end, we introduce GRAph-Structured Pyramidal (GRASP) WSI representation and use two different datasets, Ovarian Carcinoma and Bladder Cancer, to evaluate the model. Our main novelty and contributions include:
@@
    Introducing GRASP to capture pyramidal information contained in WSIs, as the first multi-magnification graph-structured model which integrates the concept of zoom-in as well as zoom-out into its structure.
@@
    Despite other multi-magnification approaches, GRASP benefits from a structural convergence instead of traditional pooling layers to capture inter-magnification information.
@@
    We also publish two large datasets of rare cancers containing 948 graphs and WSIs for Ovarian Carcinoma and 262 graphs for Bladder Cancer. The code and the datasets are publicly available at *anonymized*.
@@ 2.1 Patch-Level Encoding
With recent progress in deep learning, deep features, i.e. high-level embeddings from a deep network, have advanced past handcrafted features and are considered the most robust sources for image representation. Unfortunately, pre-trained networks such as DenseNet [14] or ResNet [12] draw their features from millions of non-medical and non-histopathological images, where they cannot necessarily produce high-level embeddings for complex images, especially rare cancers [27, 22, 6]. In this context, the use of Variational Autoencoders (VAEs) has been evaluated in [3], where the authors show that DenseNet pre-trained on ImageNet performs better for extracting semantic features from WSIs than VAEs. However, KimiaNet [22] was developed by employing the architecture of DenseNet with four dense blocks trained on a large set of pseudo-annotated histopathology images (patches) outperforming its pre-trained version on ImageNet.
@@ 2.2 Weak Supervision in Gigapixel WSIs
MIL Approaches: Several domains of deep learning have been explored in an attempt to effectively address the task of classification in digital pathology. Models such as AB-MIL [15], CLAM [18], and Trans-MIL [25] have utilized MIL with promising results. Such approaches have generally focused only on instance-level feature extraction and have not yet explored modeling global, long-range interactions within and across different magnifications.
Graph-based Approaches: To incorporate contextual information and long-range interactions, models such as PatchGCN [5] and DGCN [29] have been designed with a graph structure that can capture and learn context-aware features from interactions across the WSI. These models represent WSIs as graphs where the nodes are usually embeddings and edges are defined based on clustering or neighborhood node similarity, which in turn adds new hyperparameters and increases inference time. The similarity between nodes can be measured in terms of spatial or latent space, leading to the construction of different graphs for each WSI.
Multi-Magnification Approaches: Multiple efforts have been made to incorporate multi-magnification information in the context of gigapixel histopathology subtyping tasks. Models such as ZoomMIL [26], CSMIL [7], $H^{2}$-MIL [13], and DSMIL [17] address this by aggregating contextual tissue information using features from multiple magnifications in WSIs. DSMIL concatenates embeddings from different magnifications together by duplicating lower-magnifications features, making the model biased toward lower magnifications and unable to look into inter-magnification information. On the other hand, ZoomMIL aggregates information from $5x$ to $20x$ in a fixed hierarchy with no interaction in the opposite direction. Chen et. al explore this in the context of vision transformers with their Hierarchical Image Pyramid Transformer (HIPT) [4]. Their architecture incorporates regions of size $256\times 256$ and $4096\times 4096$ pixels to leverage the natural hierarchical structure of WSIs. The aforementioned methods result in millions of parameters, which leads to highly complex and less interpretable models. $H^{2}$-MIL also adopts a graph-based approach, where it pools the nodes in each magnification using an Iterative Hierarchical Pooling module. Our proposed model, on the other hand, is designed to dynamically aggregate information within and across different magnifications without using traditional pooling layers in its inter-magnification interactions.
@@ 3 Method
Figure 2: Overview of our workflow beginning with WSIs and outputting slide-level subtype predictions. a) shows the WSI being tiled into patches of varying magnification which are then embedded and assembled into a hierarchical graph. In b), graph representations are fed into a three-layer GCN and subsequently, a two-layer MLP to predict graph-level (slide-level) subtypes. As shown in the message passing steps in b), nodes in the first GCN layer interact with their immediate neighbors; those in the second GCN layer can interact with their second neighbors; and nodes in the final GCN layer can interact with all nodes in the graph. Then, the inter-magnification convergence causes the nodes within each magnification to converge, which is an intrinsic property of the architecture. In the end, the three converged nodes are passed through an average readout module. This dynamic helps the model to look for important messages in the entire graph, and if a node contains important information, it will be broadcast to all other nodes in the graph. The output of the GCN layers is then averaged by the readout module and passed to the FC layers. (For the sake of illustration, $m=4$ is used to show the structure of GRASP).
@@ 3.1 Problem Formulation
For a given WSI, $W_{r}\in\mathbb{R}^{N\times M\times 3}$ with label $\mathcal{Y}$, three sets of $m$ patches, $\{p_{i}\in\mathbb{R}^{n\times n\times 3}:\forall i\in[1,...,m]\}$, $\{p^{\prime}_{i}\in\mathbb{R}^{n\times n\times 3}:\forall i\in[1,...,m]\}$, and $\{p^{\prime\prime}_{i}\in\mathbb{R}^{n\times n\times 3}:\forall i\in[1,...,m]\}$ are extracted for each magnification of $\textbf{M}_{1}=5x$, $\textbf{M}_{2}=10x$, and $\textbf{M}_{3}=20x$, respectively. It is important to note that $p^{\prime\prime}_{i}$ is the high-resolution window located at the center of $p^{\prime}_{i}$, and $p^{\prime}_{i}$ is the high-resolution window located at the center of $p_{i}$. More details are provided in Patch Extraction in Supplementary Materials.
These patches provide $3m$ patches in total that are then fed into an encoder $\phi$, KimiaNet, to encode extracted patches into a lower dimension space as follows:
@@ 3.2 GRASP
The mechanism of connecting every two nodes in $\mathbb{G}_{r}$ through $\Gamma$ is premised upon an intuition of the pyramidal nature of WSIs as well as the way in which a conventional light microscope works when one switches from one magnification to another. When using a microscope, increasing magnification preserves the size of the image yet increases resolution by showing the central window of the lower magnification. This is the exact procedure we use to extract our patches in three magnifications. Therefore, for any $i$, $h_{i}$, $h^{\prime}_{i}$, and $h^{\prime\prime}_{i}$ are connected to each other via undirected edges, where this connection represents intra-magnification information contained in the features. On the other hand, for any $i$, all $h_{i}$’s contain information in $\textbf{M}_{1}$, such that they are connected to each other, forming a fully connected graph at $\textbf{M}_{1}$ magnification to represent inter-magnification information. Similarly, all $h^{\prime}_{i}$s are connected to each other and also all the $h^{\prime\prime}_{i}$s to represent inter-magnification information contained in $\textbf{M}_{2}$ and $\textbf{M}_{3}$, respectively.
Figure 2 shows a small example of such a graph for $\mathbb{G}_{r}=(V_{r},E_{r})|_{m=4}$ where blue, red, and green nodes each form a fully connected graph of size $m$; the intra-magnification relationship can also be seen via the edges between blue & red nodes as well as the red & green nodes. So far, each WSI, $W_{r}$, has been represented by graph $\mathbb{G}_{r}$ with $3m$ nodes and $\frac{(3m+1)m}{2}$ edges. These graphs are thus deployed to train the GCNs and predict the label $\mathcal{Y}$ at the output.
@@ 3.3 Graph Convolutional Layers
Following Eq. 3, we are defining $\mathcal{G}$ which includes three GCN layers. The intuition behind using three layers is that as a pathologist begins to look for a tumor in a given WSI, they use an initial magnification to find the region of interest; Once found, they consult with other magnifications to confirm their decision since the data is complicated and often confusing. Therefore, as shown in Figure 2, all nodes in the graph interact with one another in a hierarchical fashion through the GCN layers. Consequently, each node gradually gathers information from all other nodes, and therefore, if there are any important messages carried by some nodes, it is guaranteed to be broadcast to all other nodes. This dynamic and hierarchical structure imposes theoretical properties on the model which is going to be discussed here. Following the graph convolutional layer introduced in [16], the graph nodes are updated as follows:
$h_{i}^{(l+1)}=\alpha(b^{(l)}+\sum_{j\in\mathcal{N}(i)}\frac{1}{c_{ji}}h_{j}^{(l)}W^{(l)})\centering\@add@centering$ (4)
where $b^{(l)}$ is bias; $h_{i}^{(l+1)}$ is the node feature’s update of the graph at $(l+1)$-th step at $\textbf{M}_{1}$; $\mathcal{N}(i)$ is the set of neighbors of node $i$, $c_{ji}=\sqrt{|\mathcal{N}(j)||\mathcal{N}(i)|}$, where given the symmetry of the graph, all $c_{ji}$s are equal; and $\alpha(.)$ is the activation function, which is ReLU in our implementation. $h_{i}^{\prime(l+1)}$ and $h_{i}^{\prime\prime(l+1)}$ expressions follow the same logic as $h_{i}^{(l+1)}$ in terms of the parameters mentioned above.
After the last graph convolutional layer, where the inter-magnification convergence happens, the graph is passed through an average readout module to pool the three-node graph mean embedding and then it is fed into the two-layer classifier $\mathcal{C}$ to predict $\mathcal{Y}$.
3.4 Inter-Magnification Convergence
Based on the idea of distilling the information within and across magnifications to capture inter- and intra-magnification information, we now show that node features are converging in each magnification in the model. Having this, GRASP essentially encodes a graph of size $3m$ nodes to only 3 nodes. We interpret this as the model learning the information contained in the magnification through interaction with other magnifications.
Theorem 1

### v2

# GRASP: GRAph-Structured Pyramidal Whole Slide Image Representation

@@ Abstract
Cancer subtyping is one of the most challenging tasks in digital pathology, where Multiple Instance Learning (MIL) by processing gigapixel whole slide images (WSIs) has been in the spotlight of recent research. However, MIL approaches do not take advantage of inter- and intra-magnification information contained in WSIs. In this work, we present GRASP, a novel lightweight graph-structured multi-magnification framework for processing WSIs in digital pathology. Our approach is designed to dynamically emulate the pathologist’s behavior in handling WSIs and benefits from the hierarchical structure of WSIs. GRASP, which introduces a convergence-based node aggregation mechanism replacing traditional pooling mechanisms, outperforms state-of-the-art methods by a high margin in terms of balanced accuracy, while being significantly smaller than the closest-performing state-of-the-art models in terms of the number of parameters. Our results show that GRASP is dynamic in finding and consulting with different magnifications for subtyping cancers, is reliable and stable across different hyperparameters, and can generalize when using features from different backbones. The model’s behavior has been evaluated by two expert pathologists confirming the interpretability of the model’s dynamic. We also provide a theoretical foundation, along with empirical evidence, for our work, explaining how GRASP interacts with different magnifications and nodes in the graph to make predictions. We believe that the strong characteristics yet simple structure of GRASP will encourage the development of interpretable, structure-based designs for WSI representation in digital pathology. Data and code can be found in https://github.com/AIMLab-UBC/GRASP
@@ 1 Introduction
Though deep learning has revolutionized computer vision in many fields, digital pathology tasks such as cancer classification remain a complex problem in the domain. For natural images, the task usually relates to assigning a label to an image with an approximate size of 256 × 256 pixels, with the label being clearly visible and well-represented in the image. Gigapixel tissue whole-slide images (WSIs) break this assumption in digital pathology as images exhibit enormous heterogeneity and can be as large as 150,000 × 150,000 pixels. Further, labels are provided at the slide level and may be descriptive of a small region of pixels occupying a minuscule portion of the total image, or they may be descriptive of complex interactions between the substructures within the entire composition of the WSI Ehteshami Bejnordi et al. (2017); Zhang et al. (2015); Pawlowski et al. (2019).
@@
Multiple Instance Learning (MIL) has become the prominent approach to address the computational complexity of WSI; however, the majority of methods in the literature focus only on a single level of magnification, usually 20×20$\times$20 ×, due to the computational cost of including other magnifications Lu et al. (2021); Ilse et al. (2018); Schirris et al. (2022); Zheng et al. (2021); Chen et al. (2021); Shao et al. (2021); Zhou et al. (2019); Guan et al. (2022). Using this magnification, a set of patches from each WSI are extracted and used as an instance-level representation. This neither captures the biological structure of the data nor does it follow the diagnostic protocols of pathologists. That is to say, WSIs at higher magnifications reveal finer details—such as the structure of the cell nucleus and the intra/extracellular matrix—whereas lower magnifications enable the identification of larger structures like blood vessels, connective tissue, or muscle fibers. Further, these structures are inconsistent from patient to patient, slide to slide, and subtype to subtype Morovic et al. (2021). To capture this variability, pathologists generally use a variety of lenses in their inspection of a tissue sample under the microscope, switching between different magnifications as needed. They generally begin with low magnifications to identify regions of interest for making preliminary decisions before increasing magnifications to confirm or rule out diagnoses Rasoolijaberi et al. (2022).
Figure 1: A chronological overview of different WSI representation methods and their performance compared to the size of the model.
To address this challenge, several multi-magnification approaches have recently been introduced for various tasks such as cancer subtyping, survival analysis, and image retrieval. However, these models often possess millions of parameters, as briefly illustrated in Figure 1, and suffer from interpretability issues due to their modular complexity Thandiackal et al. (2022); Li et al. (2021); Riasatian et al. (2021); Chen et al. (2022b); D’Alfonso et al. (2021); Hashimoto et al. (2020). Although these models have demonstrated promise across different tasks, they are not well-suited for low-resource clinical settings, where computational resources are often limited and the infrastructure may not support large-scale computational clusters. Therefore, there is a critical need to develop approaches and models specifically designed for use in smaller clinics, where the hardware may consist of small GPUs with limited memory. These lightweight models must balance accuracy with efficiency, enabling reliable deployment on standard devices while ensuring real-time performance and ease of integration within existing clinical workflows.
In this research, we aim to further the progress of deep learning in this context by introducing a pre-defined fixed structure for a lightweight model that reduces the complexity while maintaining efficacy and interoperability. Therefore, our contribution is as follows:
@@
    Introducing GRASP to capture pyramidal information contained in WSIs, as the first lightweight multi-magnification model in computational pathology.
@@
    GRASP introduces a novel convergence-based mechanism instead of traditional pooling layers to capture intra-magnification information.
@@
    An extensive comparison with eleven state-of-the-art models across three different cancers, ranging from two to five histotypes, using two popular backbones demonstrates the generalizability of the proposed method.
@@ 2.1 Patch-Level Encoding
With recent progress in deep learning, deep features, i.e. high-level embeddings from a deep network, have advanced past handcrafted features and are considered the most robust sources for image representation. Pre-trained networks such as DenseNet Huang et al. (2017), ResNet He et al. (2016), or Swin Liu et al. (2021) draw their features from millions of non-medical and non-histopathological images, where they cannot necessarily produce high-level embeddings for complex images, especially rare cancers Wang et al. (2023); Riasatian et al. (2021); Ciga et al. (2022); Wang et al. (2022). In this context, the use of Variational Autoencoders (VAEs) has been evaluated in Chen et al. (2022a), where the authors show that DenseNet pre-trained on ImageNet performs better for extracting semantic features from WSIs than VAEs. However, domain-specific vision encoders such as KimiaNet Riasatian et al. (2021), CTransPath Wang et al. (2022), PLIP Huang et al. (2023), UNI Chen et al. (2023), Virchow Vorontsov et al. (2023), etc. were developed and trained on large sets of histopathology images (patches) outperforming pre-trained models on ImageNet across various tasks.
@@ 2.2 Weak Supervision in Gigapixel WSIs
MIL Approaches: Several domains of deep learning have been explored in an attempt to effectively address the task of classification in digital pathology. Models such as AB-MIL Ilse et al. (2018), CLAM Lu et al. (2021), and Trans-MIL Shao et al. (2021) have utilized MIL with promising results. Such approaches have generally focused only on instance-level feature extraction and have not yet explored modeling global, long-range interactions within and across different magnifications. In Waqas et al. (2024), a detailed overview of different MIL methods has been provided.
Graph-based Approaches: To incorporate contextual information and long-range interactions, models such as PatchGCN Chen et al. (2021) and DGCN Zheng et al. (2021) have been designed with a graph structure that can capture and learn context-aware features from interactions across the WSI. These models represent WSIs as graphs where the nodes are usually embeddings and edges are defined based on clustering or neighborhood node similarity, which in turn adds new hyperparameters and increases inference time. The similarity between nodes can be measured in terms of spatial or latent space, leading to the construction of different graphs for each WSI.
Multi-Magnification Approaches: Multiple efforts have been made to incorporate multi-magnification information in the context of gigapixel histopathology subtyping tasks. Models such as HiGT Guo et al. (2023), ZoomMIL Thandiackal et al. (2022), CSMIL Deng et al. (2023), $H^{2}$-MIL Hou et al. (2022), and DSMIL Li et al. (2021) address this by aggregating contextual tissue information using features from multiple magnifications in WSIs. DSMIL concatenates embeddings from different magnifications together by duplicating lower-magnification features, making the model biased toward lower magnifications and unable to look into inter-magnification information. On the other hand, ZoomMIL aggregates information from $5x$ to $20x$ in a fixed hierarchy with no interaction in the opposite direction. Chen et. al explore this in the context of vision transformers with their Hierarchical Image Pyramid Transformer (HIPT) Chen et al. (2022b). Their architecture incorporates regions of size $256\times 256$ and $4096\times 4096$ pixels to leverage the natural hierarchical structure of WSIs. $H^{2}$-MIL also adopts a graph-based approach, where it pools the nodes in each magnification using an Iterative Hierarchical Pooling module. Our proposed model, on the other hand, is designed to dynamically aggregate information within and across different magnifications without using traditional pooling layers in its intra-magnification interactions. It also employs a similar mechanism to zoom-in and zoom-out through its inter-magnification interactions, from lower to higher magnifications and vice versa.
@@ 3 Method
Figure 2: Overview of our workflow beginning with WSIs and outputting slide-level subtype predictions. a) shows the WSI being tiled into patches of varying magnification which are then embedded and assembled into a hierarchical graph. In b), graph representations are fed into a three-layer GCN Kipf & Welling (2016) and subsequently, a two-layer MLP to predict graph-level (slide-level) subtypes. As shown in the message passing steps in b), nodes in the first GCN layer interact with their immediate neighbors; those in the second GCN layer can interact with their second neighbors; and nodes in the final GCN layer can interact with all nodes in the graph. Then, the inter-magnification convergence causes the nodes within each magnification to converge, which is an intrinsic property of the architecture. In the end, the three converged nodes are passed through an average readout module. This dynamic helps the model to look for important messages in the entire graph, and if a node contains important information, it will be broadcast to all other nodes in the graph. The output of the GCN layers is then averaged by the readout module and passed to the FC layers. (For the sake of illustration, $m=4$ is used to show the structure of GRASP).
@@ 3.1 Problem Formulation
For a given WSI, $W_{r}\in\mathbb{R}^{N\times M\times 3}$ with label $\mathcal{Y}$, three sets of $m$ patches, $\{p_{i}\in\mathbb{R}^{n\times n\times 3}:\forall i\in[1,...,m]\}$, $\{p^{\prime}_{i}\in\mathbb{R}^{n\times n\times 3}:\forall i\in[1,...,m]\}$, and $\{p^{\prime\prime}_{i}\in\mathbb{R}^{n\times n\times 3}:\forall i\in[1,...,m]\}$ are extracted for each magnification of $\textbf{M}_{1}=5x$, $\textbf{M}_{2}=10x$, and $\textbf{M}_{3}=20x$, respectively. It is important to note that $p^{\prime\prime}_{i}$ is the high-resolution window located at the center of $p^{\prime}_{i}$, and $p^{\prime}_{i}$ is the high-resolution window located at the center of $p_{i}$.  These patches provide $3m$ patches in total that are then fed into an encoder $\phi$ to encode extracted patches into a lower dimension space as follows:
@@ 3.2 GRASP
The mechanism of connecting every two nodes in $\mathbb{G}_{r}$ through $\Gamma$ is premised upon an intuition of the pyramidal nature of WSIs as well as the way in which a conventional light microscope works when one switches from one magnification to another. When using a microscope, increasing magnification preserves the size of the image yet increases resolution by showing the central window of the lower magnification. This is the exact procedure we use to extract our patches in three magnifications. Therefore, for any $i$, $h_{i}$, $h^{\prime}_{i}$, and $h^{\prime\prime}_{i}$ are connected to each other via undirected edges, where this connection represents inter-magnification information contained in the features. On the other hand, for any $i$, all $h_{i}$’s contain information in $\textbf{M}_{1}$, such that they are connected to each other, forming a fully connected graph at $\textbf{M}_{1}$ magnification to represent intra-magnification information. Similarly, all $h^{\prime}_{i}$s are connected to each other and also all the $h^{\prime\prime}_{i}$s to represent intra-magnification information contained in $\textbf{M}_{2}$ and $\textbf{M}_{3}$, respectively.
Figure 2 shows a small example of such a graph for $\mathbb{G}_{r}=(V_{r},E_{r})|_{m=4}$ where blue, red, and green nodes each form a fully connected graph of size $m$; the inter-magnification relationship can also be seen via the edges between blue & red nodes as well as the red & green nodes. So far, each WSI, $W_{r}$, has been represented by a fixed graph $\mathbb{G}_{r}$ with $3m$ nodes and $\frac{(3m+1)m}{2}$ edges. These graphs are thus deployed to train the GCNs and predict the label $\mathcal{Y}$ at the output.
@@ 3.3 Graph Convolutional Layers
Following Eq. 3, we are defining $\mathcal{G}$ which includes three GCN layers. The intuition behind using three layers is that as a pathologist begins to look for a tumor in a given WSI, they use an initial magnification to find the region of interest; Once found, they consult with other magnifications, which may require zooming in and out back and forth, to confirm their final decision. Therefore, as shown in Figure 2, all nodes in the graph interact with one another in a hierarchical fashion through the GCN layers. Consequently, each node gradually gathers information from all other nodes, and therefore, if there are any important messages carried by some nodes, it is guaranteed to be broadcast to all other nodes, which is the equivalent of zoom-in and zoom-out mechanism. This dynamic and hierarchical structure imposes theoretical properties on the model which is going to be discussed here. Following the graph convolutional layer introduced in Kipf & Welling (2016), the graph nodes are updated as follows:
$h_{i}^{(l+1)}=\alpha(b^{(l)}+\sum_{j\in\mathcal{N}(i)}\frac{1}{c_{ji}}h_{j}^{(l)}W^{(l)}),\centering\@add@centering$ (4)
where $b^{(l)}$ is bias; $h_{i}^{(l+1)}$ is the node feature’s update of the graph at $(l+1)$-th step at $\textbf{M}_{1}$; $\mathcal{N}(i)$ is the set of neighbors of node $i$, $c_{ji}=\sqrt{|\mathcal{N}(j)||\mathcal{N}(i)|}$, where given the symmetry of the graph, all $c_{ji}$s are equal; and $\alpha(.)$ is the activation function, which is ReLU in our implementation. $h_{i}^{\prime(l+1)}$ and $h_{i}^{\prime\prime(l+1)}$ expressions follow the same logic as $h_{i}^{(l+1)}$ in terms of the parameters mentioned above.  After the last graph convolutional layer, where the intra-magnification convergence happens, the graph is passed through an average readout module to pool the three-node graph mean embedding and then is fed into the two-layer classifier $\mathcal{C}$ to predict $\mathcal{Y}$.
3.4 Intra-Magnification Convergence
Based on the idea of capturing the information within and across magnifications, we now show that node features converge in each magnification in the graph to one node Having this, GRASP essentially encodes a graph of size $3m$ nodes to only 3 nodes. We interpret this as the model learning the information contained in the magnification through interaction with other magnifications without a need for traditional pooling layers.
Theorem 1.